# WaveletDiff: Multilevel Wavelet Diffusion For Time Series Generation

### v1

@@ 5 Conclusion
We provide comprehensive implementation details to ensure reproducibility of our results. Model architecture specifications, hyperparameters, and training configurations are detailed in Section 3.2 and Appendix B.4. Dataset descriptions and preprocessing steps are provided in Section 4.1 and Appendix B.1. Evaluation metrics and experimental protocols are specified in Section 4.1 and Appendix B.2. Wavelet selection criteria and mother wavelet analysis are documented in Appendix B.3. Computational requirements and training procedures are outlined in Appendix B.5. All experimental settings, including cross-level attention mechanisms, energy preservation constraints, and ablation study configurations, are thoroughly described in the main paper and corresponding appendix sections to enable complete reproduction of our findings. Our code is available at https://anonymous.4open.science/r/WaveletDiff-27E9/.
@@ Appendix A Mother Wavelet Families
Different wavelet families provide distinct characteristics of multi-scale decompositions through their specific filter coefficients $\{h_{k}\}$ and $\{g_{k}\}$ in the two-scale relations (Equation 1). In our experiments, we used five representative wavelet families, each satisfying different relationships between their order $p$ and filter length $F$. The vanishing moment properties and definitions presented here follow the PyWavelets framework implementation Lee et al. (2019b).
@@ B.2 Metrics
Discriminative Score. The discriminative score captures how difficult it is for a classifier to distinguish between real and generated samples. The score is measured by $|\text{acc}-0.5|$, where acc is the classification accuracy. A score close to 0 indicates that real and generated samples are indistinguishable to the classifier, while a score close to $0.5$ indicates they are very different. We follow the setup of TimeGAN Yoon et al. (2019) using a $2$-layer GRU-based neural network as the classifier, trained with binary cross-entropy loss to distinguish between real (label=1) and synthetic (label=0) sequences.
Predictive Score. The predictive score captures how useful generated samples are for the forecasting task on real data. The score is measured by the mean absolute error (MAE) between predicted values and ground-truth values on test data. We follow TimeGAN Yoon et al. (2019) using a $2$-layer GRU-based sequence predictor trained on synthetic data to predict the next time step features, evaluated on real sequences. Lower MAE values indicate better predictive utility of the generated samples.
Context-FID. Paul et al. (2022) The Fréchet Inception Distance (FID) measures the distance between two multivariate Gaussian distributions, i.e.,
@@
where $\mu_{X},\mu_{Y}$ are the means and $\Sigma_{X},\Sigma_{Y}$ are the covariance matrices of the two distributions. Context-FID adapts this to time series by replacing the Inception-v3 features with time series features. We follow Diffusion-TS Yuan & Qiao (2024) using TS2Vec Yue et al. (2022) representations as the features. We extract embeddings from both real and generated sequences using a trained TS2Vec encoder, then compute FID in the embedding space. Lower Context-FID values indicate better distributional similarity.
@@
where $\rho^{real}_{i,j}$ and $\rho^{generated}_{i,j}$ are the correlation coefficients between features $i$ and $j$ for real and generated data, respectively. Note that we follow the Diffusion-TS Yuan & Qiao (2024) setup using the factor $\frac{1}{10}$, although $\frac{1}{D^{2}}$ could provide better normalization across different feature dimensions. The former choice of normalization ensures direct comparability with prior work.
@@ B.4 Training Configuration Details
For optimization, we use the AdamW optimizer with initial learning rate $2\times 10^{-4}$ and weight decay $1\times 10^{-5}$. We employ a one-cycle learning rate schedule Smith & Topin (2018) with cosine annealing strategy. The learning rate follows a two-phase schedule: Linear warm-up for the first $30\%$ of training, then cosine annealing for the remaining $70\%$:
@@ Appendix D Diffusion Model Reproducibility Analysis
Inspired by recent work on reproducibility in diffusion models for images Zhang et al. (2024b); Li et al. (2024); Kadkhodaie et al. (2024), we examine whether this phenomenon extends to time series generation. Specifically, we train pairs of models with slightly different architectures or configurations, then generate samples from the same fixed Gaussian noise input using deterministic DDIM sampling. For each pair of generated sequences $(x_{1},x_{2})$, we compute their similarity using dynamic time warping (DTW) distance. To quantify reproducibility, we follow Zhang et al. (2024b) and define the RP score as

### v2

@@ 5 Conclusion
We provide comprehensive implementation details to ensure reproducibility of our results. Model architecture specifications, hyperparameters, and training configurations are detailed in Section 3.2 and Appendix B.4. Dataset descriptions and preprocessing steps are provided in Section 4.1 and Appendix B.1. Evaluation metrics and experimental protocols are specified in Section 4.1 and Appendix B.2. Wavelet selection criteria and mother wavelet analysis are documented in Appendix B.3. Computational requirements and training procedures are outlined in Appendix B.5. All experimental settings, including cross-level attention mechanisms, energy preservation constraints, and ablation study configurations, are thoroughly described in the main paper and corresponding appendix sections to enable complete reproduction of our findings.
@@ Appendix A Mother Wavelet Families
Different wavelet families provide distinct characteristics of multi-scale decompositions through their specific filter coefficients $\{h_{k}\}$ and $\{g_{k}\}$ in the two-scale relations (Equation 1). In our experiments, we used five representative wavelet families, each satisfying different relationships between their order $p$ and filter length $F$. The vanishing moment properties and definitions presented here follow the PyWavelets framework implementation lee2019pywavelets.
@@ B.2 Metrics
Discriminative Score. The discriminative score captures how difficult it is for a classifier to distinguish between real and generated samples. The score is measured by $|\text{acc}-0.5|$, where acc is the classification accuracy. A score close to 0 indicates that real and generated samples are indistinguishable to the classifier, while a score close to $0.5$ indicates they are very different. We follow the setup of TimeGAN yoon2019time using a $2$-layer GRU-based neural network as the classifier, trained with binary cross-entropy loss to distinguish between real (label=1) and synthetic (label=0) sequences.
Predictive Score. The predictive score captures how useful generated samples are for the forecasting task on real data. The score is measured by the mean absolute error (MAE) between predicted values and ground-truth values on test data. We follow TimeGAN yoon2019time using a $2$-layer GRU-based sequence predictor trained on synthetic data to predict the next time step features, evaluated on real sequences. Lower MAE values indicate better predictive utility of the generated samples.
Context-FID. paul2022psaganprogressiveselfattention The Fréchet Inception Distance (FID) measures the distance between two multivariate Gaussian distributions, i.e.,
@@
where $\mu_{X},\mu_{Y}$ are the means and $\Sigma_{X},\Sigma_{Y}$ are the covariance matrices of the two distributions. Context-FID adapts this to time series by replacing the Inception-v3 features with time series features. We follow Diffusion-TS yuan2024diffusionts using TS2Vec yue2022ts2vecuniversalrepresentationtime representations as the features. We extract embeddings from both real and generated sequences using a trained TS2Vec encoder, then compute FID in the embedding space. Lower Context-FID values indicate better distributional similarity.
@@
where $\rho^{real}_{i,j}$ and $\rho^{generated}_{i,j}$ are the correlation coefficients between features $i$ and $j$ for real and generated data, respectively. Note that we follow the Diffusion-TS yuan2024diffusionts setup using the factor $\frac{1}{10}$, although $\frac{1}{D^{2}}$ could provide better normalization across different feature dimensions. The former choice of normalization ensures direct comparability with prior work.
@@ B.4 Training Configuration Details
For optimization, we use the AdamW optimizer with initial learning rate $2\times 10^{-4}$ and weight decay $1\times 10^{-5}$. We employ a one-cycle learning rate schedule smith2018superconvergencefasttrainingneural with cosine annealing strategy. The learning rate follows a two-phase schedule: Linear warm-up for the first $30\%$ of training, then cosine annealing for the remaining $70\%$:
@@ Appendix D Diffusion Model Reproducibility Analysis
Inspired by recent work on reproducibility in diffusion models for images pmlr-v235-zhang24cn; li2024understanding; kadkhodaie2024generalization, we examine whether this phenomenon extends to time series generation. Specifically, we train pairs of models with slightly different architectures or configurations, then generate samples from the same fixed Gaussian noise input using deterministic DDIM sampling. For each pair of generated sequences $(x_{1},x_{2})$, we compute their similarity using dynamic time warping (DTW) distance. To quantify reproducibility, we follow pmlr-v235-zhang24cn and define the RP score as